At what angle of polarizer and analyzer will intensity is minimum and maximum

Grade:12th pass

1 Answers

Eshan
5 years ago
After passing through polarizer, the light is polarized in its polarizing direction. When the analyzer is at an angle\thetato the direction of polarization, the incident intensityI_0is reduced to theI_0 cos^2\theta. Hence the intensity is maximum for\theta=0^{\circ}and minimum for\theta=90^{\circ}